OpenStudy (anonymous):

how do i simplify -7x-5y=1.5 -5x-4y=4.5

OpenStudy (anonymous):

-35x-25y=1.5*5 and -35x-28y=4.5*7 1.5*5+25y=4.5*7+28y 7.5-31.5=3y -24 = 3y y = -8 and x = -5.5